x(dy)/(dx)=x^(2) + 5y

1
Expert's answer
2022-04-07T17:26:45-0400

"x\frac{dy}{dx}=x^2+5y"

"y’-5\frac{y}{x}=x" (1)

Let’s solve the following equation 

"y’-5\frac{y}{x}=0"

"\frac {dy}{y}=5\frac{dx}{x}"

"y=C(x)x^{5}"

To find solution of the equation (1) we should think C is a function of x

"y'=C'x^{5}+5Cx^{4}"

Substitution y and y’ into (1) gives

"C'x^{5}+5Cx^{4} -5\frac{Cx^{5}}{x}=x"

"C'=x^{-4}"

"C=\frac{1}{-3x^3}+C_1"

"y=(\frac{1}{-3x^3}+C_1)x^5=-\frac13x^2+C_1x^5"

Answer: "y=(\frac{1}{-3x^3}+C_1)x^5=-\frac13x^2+C_1x^5" .